# The cool and the cruel: separating hard parts of LWE secrets

### v1

@@ Abstract
Sparse binary LWE secrets are under consideration for standardization for Homomorphic Encryption and its applications to private computation [18]. Known attacks on sparse binary LWE secrets include the sparse dual attack [4] and the hybrid sparse dual-meet in the middle attack [17], which requires significant memory. In this paper, we provide a new statistical attack with low memory requirement. The attack relies on some initial lattice reduction. The key observation is that, after lattice reduction is applied to the rows of a q-ary-like embedded random matrix $\mathbf{A}$, the entries with high variance are concentrated in the early columns of the extracted matrix. This allows us to separate out the “hard part” of the LWE secret. We can first solve the sub-problem of finding the “cruel” bits of the secret in the early columns, and then find the remaining “cool” bits in linear time. We use statistical techniques to distinguish distributions to identify both the cruel and the cool bits of the secret. We provide concrete attack timings for recovering secrets in dimensions $n=256$, $512$, and $768$. For the lattice reduction stage, we leverage recent improvements in lattice reduction (flatter [32]) applied in parallel. We also apply our new attack in the RLWE setting for $2$-power cyclotomic rings, showing that these RLWE instances are much more vulnerable to this attack than LWE.
@@ 1 Introduction
produces a reduced matrix $\mathbf{A}^{\prime}$ with a non-uniform distribution of variances over its columns, see Section 3 for details. Specifically, the variance of the first $n_{u}$ columns of $\mathbf{A}^{\prime}$ is unaffected by lattice reduction. We call the bits of the secret in the first $n_{u}$ columns “cruel" (unreduced). The other $n_{r}=n-n_{u}$ columns of the matrix have significantly reduced variance, and we call the secret bits corresponding to these columns the “cool" bits. The number of unreduced and reduced columns of $\mathbf{A}^{\prime}$, $n_{u}$ and $n_{r}$, depends on the overall lattice reduction quality and the penalty parameter $\omega$, which controls a trade-off between reduction and the error variance after reduction.
@@
Outline of the paper. Section 2 describes related work. Section 3 presents our attack in detail. Section 4 explains the lattice reduction step for processing LWE samples to use in the attack. Section 5 contains the statistical analysis of the attack, including an estimate for number of samples needed. Section 6 presents our concrete secret recovery results and displays performance of other attacks. Section 7 adapts the attack to the RLWE setting and shows the improvement there. See [redacted link] for the Python code for running the attack.
@@ 2 Related Work
Our work. In contrast to (but not necessarily incompatible with) work in hybrid attacks, our attack performs the lattice reduction on a full-sized lattice using different subsets of available LWE samples. Our lattice is endowed with a scaling (or penalty) parameter $\omega$. The scaling is similar to the scaling used in [9, 4], and its utility is discussed in Section 4. Our key insight, the non-uniformity over coordinates in the reduced basis, naturally splits the problem into two parts (in this case the “cruel” and “cool” bits), reminiscent of a hybrid attack.
@@ 3 The Attack
Attack Part 3: Secret Guessing. Next, we leverage statistical properties of the cool and cruel regions to guess the secret. Figure 2 shows a histogram of the distribution of the residuals $\mathbf{a}\cdot\mathbf{s}-b\mod q$ (in blue) and $\mathbf{a}\cdot\mathbf{s}^{*}-b\mod q$ (in orange), where $\mathbf{s}^{*}$ shares the first $n_{u}$ with $\mathbf{s}$ and has all other $n_{r}$ elements set to 00. Green corresponds to a random guess for the secret with equal hamming weight. The orange and the green distributions are distinct, enabling us to detect the correctness of the first $n_{u}$ bits in the secret without knowing the others. Given the size of the error on the reduced LWE problem, so far we have not found a more efficient attack for guessing the first $n_{u}$ secret bits than brute force. One could think of using one of the hybrid attack strategies outlined in the related work on the cruel part, but we leave this for future work.
Figure 2: Histogram of 4 million samples of the residual $(\mathbf{a}\cdot\mathbf{s^{*}}-b)\mod q$ (centered and normalized) for different secret guesses $\mathbf{s^{*}}$ ($n=512$, $\log_{2}q=41$, $h=20$). The three histograms correspond to three different guesses. Blue corresponds to the true secret, orange to the secret candidate that has only the first $n_{u}=75$ bits correct and green to a random secret with $h=20$. The numbers in the legend correspond to the sample standard deviation relative to the uniform standard deviation $\frac{q}{\sqrt{12}}$.
@@ 4.1 Our Reduction Methods
Early termination. To increase the number of reduced LWE matrices produced from a single lattice reduction run, we leverage an early termination strategy. When the standard deviation reduction ratio $\rho$ reaches some threshold $\tau$, we export the reduced matrix, then sub-sample a new $(m+n)\times(m+n)$ q-ary matrix from the initial set of $n$ LWE samples and start again. This better leverages available compute resources.
@@ 4.2 Final Datasets
We apply the interleaved reduction strategy described above on the $4$ different parameter sets. The number of cruel bits, $n_{u}$, is strongly correlated with the reduction factor $\rho$: better reduction (lower $\rho$) yields fewer cruel bits. Because brute-forcing the $n_{u}$ bits is the bottleneck for our attack, we optimize the lattice reduction configuration to maximize reduction. Reduction is trivially parallelizable, since each matrix can be processed on a different core, but is slower for larger dimensions $n$ and moduli $q$.
@@ 5.1 Testing brute-forced guesses of the cruel bits
  • $\sigma_{r}=0.19\sigma_{u}$ for $n=768$ and $\log_{2}q=35$ ($\rho=0.710$ and $n_{u}=373$).
When all cruel bits are correctly predicted, and if $h_{r}$ is the number of ones in the $n_{r}$ cool bits of the secret (a known value since the Hamming weight of the secret is known), the variance of $x$ is $\sigma^{2}_{\mathcal{G}_{q}}(x)=F_{q}(h_{r}\sigma_{r}^{2}+\sigma_{e}^{2})$, with $F_{q}(v)$ the variance modulo $q$ of a centered Gaussian with variance $v$ and a modular operation. We assume independence in the reduced samples and therefore a Gaussian distribution of $x$ before the modulo operation, so the Central Limit Theorem (CLT) holds.
@@
where $G^{-1}$ is the percent point function of $\mathcal{N}(0,1)$ and $\sigma_{\mathcal{U}_{q}}^{2}$ (resp. $\sigma^{2}_{\mathcal{U}_{q}^{2}}$) is the distribution variance of $x$ (resp. $x^{2}$) under the null hypothesis of uniformity. $\alpha$ is the type-I error, i.e. the probability of rejecting a true null hypothesis. Since we are brute-forcing a large number of possible solutions, the type-I error, $\alpha$ should be small (i.e. $\alpha<$ the inverse of the number of secret candidates).
Given $\alpha$ and $\beta$ type-II error (the probability of failing to reject a false null hypothesis), we can estimate the minimum number of samples needed for our attack to succeed with probability $\approx 1-\beta$ using CLT (see 0.A.2 for more details):
@@
Table 5: Number of samples $M$ needed in the statistical test for type-I error $\alpha=2^{-128}$ and type-II error $\beta=10^{-5}$ for LWE settings specified by $n$, $\log_{2}q$, $n_{u}$, and secret Hamming weight $h$. The Worst Case number of samples corresponds to the case where all the 1s in the secret are in the reduced region $h_{r}=h$, whereas the Average case assumes the average number of 1s in that region, which is the fraction $\frac{n_{r}}{n}h$
@@
Table 6: Concrete attacks on different secrets: $n=$ dimension, $\log_{2}q$ is the modulus size, $h=$ Hamming weight of the secret. $h_{u}$ indicates the number of cruel bits which are $1$ and cumulative % measure how many secrets with hamming weight $h$ have $h_{u}$ bits and $h_{u}$ bits or fewer, respectively. We run all attacks on one GPU to be comparable, but also demonstrate scaling in the toughest setting to 20 GPUs.
@@ 6 Recovery Results
All times reported in Table 6 include the compilation time during the first run over a secret, which is of the order of 10 seconds, and loading data, also of the order of 10 seconds. For some secrets, we report multiple timings to illustrate that the process is somewhat noisy, mostly depending on where in the enumeration the secret happens to be. For the settings with higher hamming weights ($h=60$), we use 200k samples for the brute force attack and 1M for the greedy attack. For the lower hamming weights, we use 10k-30k samples for the brute force attack and 200k for the greedy part.
[…]
Table 7: Estimates of Dual Hybrid MiTM performance [8] and concrete attack performance by [19] at parameter settings explored in this work. ROP = estimated number of required operations for attack, Memory = estimated memory required for MiTM hash table, Repeats = ##\## of times attack must run to succeed with probability $0.99999$.
LWE setting ($n,q,h$) Hybrid MiTM Dual [8] Ducas et al [19]
ROP Memory Repeats Time (sec)
$(256,12,12)$ $2^{50.2}$ $3.3$TB $2^{13.6}$ 200
$(512,28,12)$ $2^{48.3}$ $2.05$TB $2^{10.4}$
$(512,41,60)$ $2^{50.4}$ $10.8$TB $2^{7.1}$
$(768,35,12)$ $2^{48.8}$ $725$GB $2^{10.6}$
Table 7 reports these results. For the estimator [8], we report the estimated number of operations, the memory required, and the number of attack repeats required to achieve a $1-10^{-5}$ probability of success (to match our $\alpha$, $\beta$ levels in §5). We convert the predicted memory into bytes by multiplying estimator output (number of integers to be stored) by number of bits required to store this integer ($32$ is suggested). We note that the ROP metric reported by the estimator is a crude estimate for runtime, since it refers to the number of comparisons or operations necessary for the attack to run, which does not easily translate to a fixed time metric. For [19], we report time to successful attack, if it succeeds. A ‘-’ in the column indicates the attack ran for $\geq 72$ hours (on the same compute hardware as our attack) without finishing.
@@
Our attack can also be applied in the 2-power cyclotomic Ring-LWE setting. Ring-LWE (RLWE) is a special case of LWE where the LWE samples can be represented more compactly as polynomials in a polynomial ring. Not all instances of Ring-LWE are hard, as was shown in [20]. But $2$-power cyclotomic rings are not vulnerable to the attack in [20], so those are the rings standardized for use in Homomorphic Encryption [2].
Consider the $2$-power cyclotomic ring defined by the polynomial $x^{n}+1$, where $n=2^{k}$: $R_{q}=\mathbb{Z}_{q}[X]/(X^{n}+1)$. Then the RLWE samples corresponding to a polynomial RLWE sample, $(a(x),b(x)=a(x)\cdot s(x)+e(x))$ can be described via a skew-circulant matrix, $\mathbf{A}_{circ}$. If $a(x)=a_{0}+a_{1}x+...+a_{n-1}x^{n-1}$, then:
@@
When the matrix has this structure, we can rotate the “cruel bits” around (without redoing any lattice reduction) and increase our chances of recovering the secret, if there are only a small number of cruel bits in one of the rotated cruel regions. This makes the RWLE problem clearly easier than generic LWE in those cases, with respect to our attack.
After reduction, we have reduced samples, $\mathbf{A}\in\mathbb{Z}_{q}^{m\times n}$, where $\mathbf{A}=\mathbf{RA}_{\text{circ}}$ for some $\mathbf{R}\in\mathbb{Z}^{m\times n}$. We can shift-negate the reduced samples $k$ times to get new samples $(\mathbf{A}^{\leftarrow k},\mathbf{b}^{\leftarrow k})$ (see 0.A.1). The attack for RLWE consists of performing brute force on $n$ sliding windows $[k,k+n_{u}\mod{n})$ simultaneously (or any other number of windows by using a higher sliding step). The exhaustive search will cover less search space as we only need to check up to $h^{(1)}_{u}=\min_{0\leq i<n}h_{n_{u},i}(\mathbf{s})$ hamming weight where
@@
In this context, $h_{n_{u},i}(\mathbf{s})$ represents the Hamming weight of a specific segment of the secret vector $\mathbf{s}$, starting at position $i$ and of length $n_{u}$. It’s important to note that when $i=0$, this notation aligns with our previous use of $h_{u}$ to denote the Hamming weight of the first $n_{u}$ elements of $\mathbf{s}$. Similarly, when $n_{r}=n-n_{u}$, $h_{n_{r},n_{u}}(\mathbf{s})$ aligns with our previous use of $h_{r}$ to denote the Hamming weight of the last $n_{r}=n-n_{u}$ elements of $\mathbf{s}$.
To evaluate the speed up of this attack compared to the LWE case, we estimate the cost of the brute force component while assuming equal cost of the lattice reduction for LWE and RLWE. Let $c\cdot M(h_{r}(s),\alpha,\beta)$ represent the cost of checking a secret candidate using $M(h_{r}(s),\alpha,\beta)$ samples, where $c$ is a constant factor that accounts for the computational resources required for each verification. Let $T_{\text{guess}}^{\text{LWE}}$ be the average time complexity of the exhaustive search for LWE, assuming we sweep through the secret candidates in ascending Hamming weight order:
@@
By using the same number of samples $M=M(h_{r}(s),\alpha,\beta)$ in the attack for all secret candidates, we have
$T_{\text{guess}}^{\text{LWE}}=c\cdot M\cdot\sum_{k=0}^{h}\binom{n_{u}}{k}\mathbb{P}(h_{u}(s)\geq k)$
In the RLWE case, we have $\mathbb{P}(h^{(1)}_{u}\leq\frac{hn_{u}}{n})=1$ so that
$T_{\text{guess}}^{\text{RLWE}}=n\cdot c\cdot M\cdot\sum_{k=0}^{\lfloor\frac{hn_{u}}{n}\rfloor}\binom{n_{u}}{k}\mathbb{P}(h^{(1)}_{u}\geq k)$
Figure 3: Exhaustive search cost ratio for LWE versus RLWE assuming a fixed $10\%$ binary secret sparsity. The values of $n_{u}$ are obtained experimentally.
In Figure 3, we estimate the ratio of the costs above. Note that the distribution of $h_{u}(s)$ conditionally to a fixed total hamming weight $h$ follows a hyper-geometric distribution when the secret bits are from Bernoulli, while the $h^{(1)}_{u}$ distribution is computed empirically.
We conduct several experiments comparing LWE with RLWE, using identical secrets for comparison. The results of these experiments, for parameters $n=256$ and $n=512$, are presented in Table 8 and Table 9. Each table row corresponds to a distinct secret $s$, and the columns detail the actual and estimated time costs for the LWE and RLWE attacks, along with the count of cruel bits in the secret. In the tables, Actual Time is the empirically measured duration of the attack, while the Estimated Time is estimated using an approximation $T_{\text{guess}}^{\text{LWE}}=b+a\sum_{k=0}^{h_{u}(s)}\binom{n_{u}}{k}$. This estimate provides an idea of the time required to attack a secret when $h_{u}$ is very large, without actually running the computation.
Table 8: Attack times for parameters: $n=256,h=12,\log q=12,n_{u}=145$ run on 20 randomly sampled secrets. Actual time is the empirical time of the attack. Estimated time is computed via the formula: $T_{\text{guess}}^{\text{LWE}}=b+a\sum_{k=0}^{h_{u}(s)}\binom{n_{u}}{k}$ and $T_{\text{guess}}^{\text{RLWE}}=n(b+a\sum_{k=0}^{h^{(1)}_{u}}\binom{n_{u}}{k})$ for RLWE. $h_{u}(s)$, as defined in Equation 16, is the number of the cruel bits of the secret, and $h^{(1)}_{u}$ is the minimum hamming weight over all windows of length $n_{u}$. We use $56k$ samples for Cruel bit recovery and $200k$ for Algorithm 1.
@@
4 $4.60\times 10^{2}$ $4.12\times 10^{2}$ 5 $9.77\times 10^{3}$ $1.02\times 10^{4}$ 4
5 $4.60\times 10^{2}$ $2.61\times 10^{2}$ 5 $5.97\times 10^{3}$ $4.58\times 10^{3}$ 3
6 $4.60\times 10^{2}$ $3.61\times 10^{2}$ 5 $9.77\times 10^{3}$ $1.19\times 10^{4}$ 4
7 $1.04\times 10^{4}$ 6 $5.97\times 10^{3}$ $6.00\times 10^{3}$ 3
8 $2.09\times 10^{5}$ 7 $1.18\times 10^{5}$ $1.76\times 10^{5}$ 5
@@
12 $2.09\times 10^{5}$ 7 $5.97\times 10^{3}$ $5.71\times 10^{3}$ 3
13 $2.09\times 10^{5}$ 7 $1.18\times 10^{5}$ $5.38\times 10^{4}$ 5
14 $3.66\times 10^{6}$ 8 $9.77\times 10^{3}$ $1.06\times 10^{4}$ 4

### v2

@@ Abstract
Sparse binary LWE secrets are under consideration for standardization for Homomorphic Encryption and its applications to private computation [20]. Known attacks on sparse binary LWE secrets include the sparse dual attack [5] and the hybrid sparse dual-meet in the middle attack [19], which requires significant memory. In this paper, we provide a new statistical attack with low memory requirement. The attack relies on some initial lattice reduction. The key observation is that, after lattice reduction is applied to the rows of a q-ary-like embedded random matrix $\mathbf{A}$, the entries with high variance are concentrated in the early columns of the extracted matrix. This allows us to separate out the “hard part” of the LWE secret. We can first solve the sub-problem of finding the “cruel” bits of the secret in the early columns, and then find the remaining “cool” bits in linear time. We use statistical techniques to distinguish distributions to identify both cruel and cool bits of the secret. We recover secrets in dimensions $n=256,512,768,1024$ and provide concrete attack timings. For the lattice reduction stage, we leverage recent improvements in lattice reduction (flatter [34]) applied in parallel. We also apply our new attack to RLWE with $2$-power cyclotomic rings, showing that these RLWE instances are much more vulnerable to this attack than LWE.
@@ 1 Introduction
produces a reduced matrix $\mathbf{A}^{\prime}$ with a non-uniform distribution of variances over its columns; see Section 3 for details. Specifically, the variance of the first $n_{u}$ columns of $\mathbf{A}^{\prime}$ is unaffected by lattice reduction. We call the bits of the secret in the first $n_{u}$ columns “cruel" (unreduced). The other $n_{r}=n-n_{u}$ columns of the matrix have significantly reduced variance, and we call the secret bits corresponding to these columns the “cool" bits. The number of unreduced and reduced columns of $\mathbf{A}^{\prime}$, $n_{u}$ and $n_{r}$, depends on the overall lattice reduction quality and the penalty parameter $\omega$, which controls a trade-off between reduction and the error variance after reduction.
@@
Outline of the paper. Section 2 describes related work. Section 3 presents our attack in detail. Section 4 explains the lattice reduction step for processing LWE samples to use in the attack. Section 5 contains the statistical analysis of the attack, including an estimate for number of samples needed. Section 6 presents our concrete secret recovery results and performance of comparable attacks. Section 7 adapts the attack to the RLWE setting.
@@ 2 Related Work
Our work. In contrast to (but not necessarily incompatible with) work in hybrid attacks, our attack performs partial lattice reduction on full-sized lattices arising from different subsets of available LWE samples. Our lattice reduction uses a scaling (or penalty) parameter $\omega$ similar to the one in [10, 5], and its utility is discussed in Section 4. Our key insight, the non-uniformity over coordinates in the reduced basis, naturally splits the problem into two parts (in this case the “cruel” and “cool” bits), reminiscent of a hybrid attack.
@@ 3 The Attack
Attack Part 3: Secret Guessing. Next, we leverage statistical properties of the cool and cruel regions to guess the secret. Figure 2 shows a histogram of the distribution of the residuals of reduced data $\mathbf{a}\cdot\mathbf{s}-b\mod q$ (in blue) and $\mathbf{a}\cdot\mathbf{s}^{*}-b\mod q$ (in orange), where $\mathbf{s}^{*}$ shares the first $n_{u}$ entries with $\mathbf{s}$ and has all other $n_{r}$ elements set to 00. Green corresponds to a random guess for the secret with equal hamming weight. The orange and the green distributions are distinct, enabling us to detect the correctness of the first $n_{u}$ bits in the secret without knowing the others. Given the size of the error on the reduced LWE problem, so far we have not found a more efficient attack for guessing the first $n_{u}$ secret bits than brute force. One could think of using one of the hybrid attack strategies outlined in the related work on the cruel part, but we leave this for future work.
Figure 2: Histogram of 4 million samples of the residual $(\mathbf{a}\cdot\mathbf{s^{*}}-b)\mod q$ (centered and normalized) for reduced data and different secret guesses $\mathbf{s^{*}}$ ($n=512$, $\log_{2}q=41$, $h=20$). The three histograms correspond to three different guesses. Blue corresponds to the true secret, orange to the secret candidate where the first $n_{u}=75$ entries are guessed correctly, and green to a random secret with $h=20$. The numbers in the legend correspond to sample standard deviation relative to uniform standard deviation $\frac{q}{\sqrt{12}}$.
@@ 4.1 Our Reduction Methods
Early termination. To increase the number of reduced LWE matrices produced from a single lattice reduction run, we leverage an early termination strategy. When the standard deviation reduction ratio $\rho$ stalls , we export the reduced matrix, then sub-sample a new $(m+n)\times(m+n)$ q-ary matrix from the initial set of $n$ LWE samples and start again. This better leverages available compute.
@@ 4.2 Final Datasets
We apply the interleaved reduction strategy described above on the $5$ different parameter sets. The number of cruel bits, $n_{u}$, is strongly correlated with the reduction factor $\rho$: better reduction (lower $\rho$) yields fewer cruel bits. Because brute-forcing the $n_{u}$ bits is the bottleneck for our attack, we optimize the lattice reduction configuration to maximize reduction. Reduction is trivially parallelizable, since each matrix can be processed on a different core, but is slower for larger dimensions $n$ and moduli $q$.
@@ 5.1 Testing Brute-forced Guesses of the Cruel Bits
  • $\sigma_{r}=0.19\sigma_{u}$ for $n=768$ and $\log_{2}q=35$ ($\rho=0.710$ and $n_{u}=373$)
  • $\sigma_{r}=0.14\sigma_{u}$ for $n=1024$ and $\log_{2}q=50$ ($\rho=0.703$ and $n_{u}=495$)
When all cruel bits are correctly predicted, and if $h_{r}$ is the number of ones in the $n_{r}$ cool bits of the secret (a known value since the Hamming weight of the secret is known), the variance of $x$ is $\sigma^{2}_{\mathcal{G}_{q}}(x)=F_{q}(h_{r}\sigma_{r}^{2}+\sigma_{e}^{2})$, with $F_{q}(v)$ the variance of a centered Gaussian modulo $q$ (which we call $\mathcal{G}_{q}$) with variance $v$ and a modular operation. We assume independence in the reduced samples and therefore a Gaussian distribution of $x$ before the modulo operation, so the Central Limit Theorem (CLT) holds.
@@
where $G^{-1}$ is the percent point function of $\mathcal{N}(0,1)$ and $\sigma_{\mathcal{U}_{q}}^{2}$ (resp. $\sigma^{2}_{\mathcal{U}_{q}^{2}}$) is the distribution variance of $x$ (resp. $x^{2}$) under the null hypothesis of uniformity. $\alpha$ is the false negative error, i.e. the probability of rejecting a true null hypothesis. Since we are brute-forcing a large number of possible solutions, the false negative error $\alpha$ should be small (i.e. $\alpha<$ the inverse of the number of secret candidates).
Given $\alpha$ and $\beta$ false positive error (the probability of failing to reject a false null hypothesis), we can estimate the minimum number of samples needed for our attack to succeed with probability $\approx 1-\beta$ using CLT (see 0.A.2 for more details):
@@
Table 5: Number of reduced samples $M$ needed in the statistical test for false negative error $\alpha=2^{-128}$ and false positive error $\beta=10^{-5}$ for LWE settings specified by $n$, $\log_{2}q$, $n_{u}$, and secret Hamming weight $h$. The Worst Case number of samples corresponds to the case where all the 1s in the secret are in the reduced region $h_{r}=h$, whereas the Average case assumes the average number of 1s in that region, which is the fraction $\frac{n_{r}}{n}h$. Because overall lattice reduction depends on both $n$ and $\log_{2}q$, problem difficulty, and thus $M$, does not monotonically increase with respect to only $n$.
@@
Table 6: Concrete attacks on different secrets. $n=$ dimension, $\log_{2}q=$ modulus size, $h=$ Hamming weight of the secret. $h_{u}=$ cruel bits equal to $1$ and cumulative % show the percentage of hamming weight $h$ secrets which have $h_{u}$ cruel bits (or fewer). All attacks run on 1 GPU, except the toughest settings on 20 GPUs in parallel.
@@ 6 Recovery Results
All times reported in Table 6 include the compilation time during the first run over a secret, which is of the order of 10 seconds, and loading data, also of the order of 10 seconds. For some hamming weights, we report multiple timings to illustrate that the process is somewhat noisy, mostly depending on where in the enumeration the secret happens to be. For the settings with higher hamming weights ($h=60$), we use 200k samples for the brute force attack and 1M for the greedy attack. For the lower hamming weights, we use 10k-30k samples for the brute force attack and 200k for the greedy part. Recall that samples are generated via the pre-processing step explained in Section 4.1, so the raw number of samples needed for each attack is always $4n$. Also note that attacks are run on 1 V100 GPU in general, using 20 GPUs in parallel only for the hardest cases.
Existing attacks. For completeness, we compare the performance of our attack to prior work. The hybrid dual meet-in-the-middle (MiTM) attack is most similar to ours [19], but options for comparison are limited, since no concrete performance results are provided in [19]. Some concrete performance results are given for related attacks, but  [8] only considers $n\leq 110$, $\log q=11$, while  [32] considers $n\leq 100$, $\log q=8$. Thus, we choose two routes for comparison. First, we compare against estimated cost of the hybrid dual MiTM estimate from the LWE estimator tool [9]$+$https://github.com/malb/lattice-estimator, commit 00ec72ce, and then we compare against a recent concrete attack implementation for sparse secret LWE proposed in [21]. The latter has code available$+$https://github.com/lducas/leaky-LWE-Estimator/tree/human-LWE/human-LWE and claims to be fast. It provides a MITM attack in Python for the $n=256$, $\log q=12$ setting and a modified uSVP/BDD attack for larger parameters.
Table 7: Claimed attacks for comparison from [9] and [21]
Estimates from [9] do not represent actual secret recovery, only predictions. Estimates given in terms of ROP = estimated number of required operations for attack, Repeats = ##\## of times attack must run to succeed with probability $0.99999$, Total Cost = total estimated ROP for this attack, omitting storage costs. For the code from [21], ‘-’ indicates no success after running for 3 weeks on our hardware.
LWE setting ($n,q,h$) Dual Hybrid MiTM [9] Ducas et al [21]
ROP Repeats Total Cost Time (sec)
$(256,12,12)$ $2^{52.7}$ $2^{14.5}$ $2^{67.2}$ 200
$(512,28,12)$ $2^{48.6}$ $2^{10.6}$ $2^{59.2}$
$(512,41,60)$ $2^{50.6}$ $2^{8.6}$ $2^{59.2}$
$(768,35,12)$ $2^{49.1}$ $2^{10.6}$ $2^{59.7}$
$(1024,50,17)$ $2^{50.9}$ $2^{11.1}$ $2^{61.0}$
Table 7 reports comparison results. For the estimator [9], we report the estimated number of operations and the number of attack repeats required to achieve a $1-10^{-5}$ probability of success (to match our $\alpha$, $\beta$ levels in §5). We note that the ROP metric reported by the estimator is a crude estimate for runtime, since it refers to the number of comparisons or operations necessary for the attack to run, which does not easily translate to time measurements. For example, if it means $(\mathbf{Z}/q\mathbf{Z})$ ring operations then one should multiply by the cost of a multiplication modulo $q$, naively $(\log q)^{2}$, or if it means ciphertext operations then one should multiply by the cost of polynomial multiplications modulo $q$, etc.
For the attack reported in [21], we find that it does not succeed and does not recover secrets in the dimensions $>256$ reported in this paper and in [26]. The attack ran for up to 3 weeks in all cases where dimension $>256$ (on the same hardware as our attack). No secrets were recovered and in some cases the jobs crashed after running for days/weeks. This claimed attack and code was also presented in [2], without verifying the code or the fact that it does not recover secrets in the dimensions ($>256$) we attack.
@@
Our attack can also be applied in the 2-power cyclotomic Ring-LWE setting. Ring-LWE (RLWE) is a special case of LWE where the LWE samples are represented more compactly as polynomials in a polynomial ring. Not all instances of Ring-LWE are hard, as was shown in [22]. But $2$-power cyclotomic rings are not vulnerable to the attack of [22], so those are the rings proposed for standardized use in Homomorphic Encryption [3, 12].
Consider the $2$-power cyclotomic ring defined by the polynomial $x^{n}+1$, where $n=2^{k}$: $R_{q}=\mathbb{Z}_{q}[x]/(x^{n}+1)$. Then the RLWE samples $(a(x),b(x)=a(x)\cdot s(x)+e(x))$ can be described via a skew-circulant matrix, $\mathbf{A}_{circ}$. If $a(x)=a_{0}+a_{1}x+...+a_{n-1}x^{n-1}$, then using the embedding $\mathbf{a}=(a_{0},a_{1},\dots,a_{n-1})$:
@@
Figure 3: Exhaustive search cost ratio for LWE versus RLWE as defined in Equation 13 and Equation 14 assuming a fixed $10\%$ binary secret sparsity. The values of $n_{u}$ are obtained experimentally.
When the matrix has this structure, we can rotate the “cruel bits” around (without redoing any lattice reduction) by constructing datasets composed of elements from certain indices in the matrix As Appendix Figure 4 shows, shifting increases our chances of recovering the secret, since we can rotate the cruel region essentially for free. Rotating around, we may find a rotation of the cruel region with only a few cruel (active) bits. We can then run our attack on this new, easier problem. This ability to shift the cruel region makes the RWLE problem clearly easier than generic LWE, with respect to our attack.
More formally, RLWE cruel bit shifting works as follows. After running reduction on the original RLWE polynomials $a(x)$, we have reduced samples, $\mathbf{A}\in\mathbb{Z}_{q}^{m\times n}$, where $\mathbf{A}=\mathbf{RA}_{\text{circ}}$ for some $\mathbf{R}\in\mathbb{Z}^{m\times n}$. We can shift-negate the reduced samples $k$ times to get new samples $(\mathbf{A}^{\leftarrow k},\mathbf{b}^{\leftarrow k})$ (see 0.A.1). The attack for RLWE brute forces $n$ sliding windows $[k,k+n_{u}\mod{n})$ simultaneously (or any other number of windows by using a higher sliding step). Exhaustive search only requires checking up to $h^{(1)}_{u}=\min_{0\leq i<n}h_{n_{u},i}(\mathbf{s})$ hamming weight where
@@
$h_{n_{u},i}(\mathbf{s})$ represents the Hamming weight of a specific segment of the secret vector $\mathbf{s}$, starting at position $i$ and of length $n_{u}$. When $i=0$, this notation aligns with our previous use of $h_{u}$ to denote the Hamming weight of the first $n_{u}$ elements of $\mathbf{s}$. Similarly, when $n_{r}=n-n_{u}$, $h_{n_{r},n_{u}}(\mathbf{s})$ aligns with our previous use of $h_{r}$ to denote the Hamming weight of the last $n_{r}=n-n_{u}$ elements of $\mathbf{s}$.
To evaluate the speed up of the RLWE attack, we estimate the cost of the brute force component while assuming equal cost of lattice reduction for LWE and RLWE. Let $c\cdot M(h_{r}(s),\alpha,\beta)$ represent the cost of checking a secret candidate using $M(h_{r}(s),\alpha,\beta)$ samples, where $c$ is a constant accounting for the computational cost of each verification. Let $T_{\text{guess}}^{\text{LWE}}$ be the average time complexity of the exhaustive search for LWE, assuming we sweep through secret candidates in ascending Hamming weight order:
@@
Using the same number of samples $M=M(h_{r}(s),\alpha,\beta)$ for each secret candidate:
$T_{\text{guess}}^{\text{LWE}}=c\cdot M\cdot\sum_{k=0}^{h}\binom{n_{u}}{k}\mathbb{P}(h_{u}(s)\geq k)$ (13)
Contrast this with the RLWE case, where we have $\mathbb{P}(h^{(1)}_{u}\leq\frac{hn_{u}}{n})=1$ so that:
$T_{\text{guess}}^{\text{RLWE}}=n\cdot c\cdot M\cdot\sum_{k=0}^{\lfloor\frac{hn_{u}}{n}\rfloor}\binom{n_{u}}{k}\mathbb{P}(h^{(1)}_{u}\geq k)$ (14)
In Figure 3, we estimate the ratio of these costs. The distribution of $h_{u}(s)$ with fixed total hamming weight $h$ follows a hyper-geometric distribution when the secret bits are from a Bernoulli distribution, while the $h^{(1)}_{u}$ distribution is computed empirically.
Table 8: RLWE/LWE Attack times for $n=256,h=12,\log q=12,n_{u}=145$, run on 10 randomly sampled secrets. Actual time is the empirical time of the attack. Estimated time is computed via the formula: $T_{\text{guess}}^{\text{LWE}}=b+a\sum_{k=0}^{h_{u}(s)}\binom{n_{u}}{k}$ and $T_{\text{guess}}^{\text{RLWE}}=n(b+a\sum_{k=0}^{h^{(1)}_{u}}\binom{n_{u}}{k})$ for RLWE. $h_{u}(s)=\#$ cruel bits of the secret, and $h^{(1)}_{u}$ is the minimum hamming weight over all windows of length $n_{u}$. We use $56K$ samples for cruel bit recovery and $200K$ for Algorithm 1.